# Decision Synthesis in Monetary Policy

### v1

@@ Abstract
The macroeconomy is a sophisticated dynamic system involving significant uncertainties that complicate modelling. In response, decision makers consider multiple models that provide different predictions and policy recommendations which are then synthesized into a policy decision. In this setting, we introduce and develop Bayesian predictive decision synthesis (BPDS) to formalize monetary policy decision processes. BPDS draws on recent developments in model combination and statistical decision theory that yield new opportunities in combining multiple models, emphasizing the integration of decision goals, expectations and outcomes into the model synthesis process. Our case study concerns central bank policy decisions about target interest rates with a focus on implications for multi-step macroeconomic forecasting.
@@ Introduction
Monetary policy makers are tasked with simple, but hard to achieve, objectives. A common objective is to target the future inflation rate, or other macroeconomic outcomes, using the interest rates as the policy instrument. These decisions are made based on uncertain information from many sources. In this paper, these sources are econometric models that generate predictive distributions for the macroeconomic outcomes and the policy instruments over multiple time periods. For a single model, it is straightforward to select an optimal policy instrument using decision analysis and conditional forecasting. Applying standard methods– such as Bayesian model averaging (BMA)– is one way to address the issue of model uncertainty, as routine decision analysis can then be applied to the weighted average as a single model. However, this traditional view ignores the reality that a set of models may each individually recommend very different optimal policy decisions. The question then arises as to how to synthesize this information and, potentially, exploit it in the overall final decision process. This paper addresses and answers this question.
@@
There is an extensive Bayesian econometrics literature on model combination, but discussion of the question that models are built for purposes– specific prediction and decision goals– is very sparse. Traditional BMA analysis weights models according to purely statistical model fit, and in time series explicitly and only scores 1-step ahead forecast outcomes. Extensions and alternatives have arisen to define model weightings based on aspects of past forecast performance with respect to specific forecast goals. Martin et al. (2023) survey Bayesian forecasting in economics and finance and overviews various forecast combination approaches, including some that are more explicitly concerned with goal-focused prediction (e.g. Mitchell and Hall, 2005; Geweke and Amisano, 2011; Conflitti et al., 2015; Kapetanios et al., 2015; Loaiza-Maya et al., 2021; Chernis and Webley, 2022; Aastveit et al., 2023; Bernaciak and Griffin, 2024). Lavine et al. (2021) provide additional perspectives and put many of the earlier approaches in a foundational Bayesian context, justifying model weights based on utilities in forecasting using historical model-specific “scoring” of past forecast outcomes. The underlying theoretical justifications come from Bayesian predictive synthesis (BPS) and the specific class of “mixture BPS” models (McAlinn and West, 2019, section 2.2; Johnson and West, 2022). However, while ultimate decision goals may be implicit in specific applications of model combination, they are rarely if ever taken into account in the analysis and resulting decision-making. This raises questions: A model that has fit or forecast specific outcomes well in the past may be a good bet for use in resulting decision analysis– in our settings, to define optimal decisions about values of policy instruments– but, there is no guarantee that this will be so.
@@
Our view is that models that have recommended optimal policy decisions that turn out to be “good” should be more heavily weighted in looking ahead, just as past statistical predictive performance is– and should be– generally positively weighted. The catch, of course, is defining “good”; econometric models were and are not explicitly used and scored in past policy decisions. The challenge is then to operationalize the concept of “good decision” performance. For example, a vector autoregression (VAR) model can be evaluated on forecast performance using a pseudo real-time forecasting exercise. But, it is not clear how to evaluate such a model when used to advise policy decisions. We can, however, use the VAR model at present to inform near-term decisions and explore how it would have advised on decisions in the past. Evaluations can then compare such analysis to decisions actually made by policy makers in the past (albeit recognizing that past decisions of policy makers were not necessarily correct – rather, just the outcomes of the amorphous reality of monetary policy-making).
@@
Bayesian Predictive Decision Synthesis (BPDS–Tallman and West, 2023; Tallman, 2024) addresses these questions. As part of the theoretical framework of Bayesian predictive synthesis (BPS–McAlinn and West, 2019; Johnson and West, 2022), BPDS explicitly allows and encourages scoring of models based on decision analysis performance as well as statistical predictive accuracy. In addition to reflecting historical outcomes of predictions and decisions, BPDS critically also allows for differential model weighting based on expected decision outcomes. This is a complete decision parallel to the proven use of BPS models that incorporate outcome-dependent weights that modify BMA-like mixtures to differentially favour models in different parts of the future outcome space for pure forecasting. The latter concept was introduced by Kapetanios et al. (2015), whose empirically inspired developments recognized, for example, that one model may be better at predicting inflation when inflation is high and rising, while another model may be better when inflation is low and stable. BPS defines the conceptual and theoretical Bayesian bases and a broader methodological framework for this. BPDS goes further by integrating both historical and expected decision outcomes; here we develop, extend, and exemplify BPDS in our central macroeconomic policy context.
@@
BPDS applies the broader Bayesian mixture model approach of BPS using defined utility– or “score”–functions that relate to explicit decision goals. Importantly, this allows for multiple objectives, i.e., multi-attribute decision analysis. For example, a purely predictive vector score function can allow for multiple forecast horizons (e.g. to produce inflation near a target for each of the next eight quarters) and/or multiple outcome criteria (e.g. to separately reflect inflation targeting, interest rate smoothing, and stable growth patterns over coming quarters), among others. For policy makers juggling multiple objectives, this is a key feature; it is rather distinct from conventional approaches that adopt single, scalar criteria for model weighting. For example, a forecast combination approach might choose model weights based on the $h-$step ahead predictive likelihood for a single choice of $h$, with BMA simply focused on $h=1$, whereas BPDS can address multi-steps ahead in parallel, along with scoring of outcomes of decision goals that simultaneously target several macroeconomic outcomes.
The opportunities for exploring such practically relevant questions, and the differences relative to traditional single-model and BMA-based analyses, are showcased in our empirical studies. This involves an exploratory case study using US macroeconomic data with multiple-objective score functions to define BPDS model weights, exemplifying the use of BPDS in macroeconomic forecasting and advisory decision-making.
@@
At a given time point, let $\mathbf{y}$ denote the $q-$dimensional outcome variable of interest (e.g. inflation in each of the next $q$ quarters) and $\mathbf{x}$ the vector of control/decision variables (e.g., a target profile of central bank interest/base rates over the next $q$ quarters). Each of a set of $J$ models ${\mathcal{M}}_{j}$, $j=1{:}J,$ predicts the outcome $\mathbf{y}$ via a predictive density $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ conditional on any considered decision $\mathbf{x}.$ The policy-maker responsible for ultimate decisions adopts a general BPDS approach that implies the overall conditional (on $\mathbf{x}$) predictive pdf of the form
@@
The decision dependent model probabilities $\pi_{j}(\mathbf{x})$ can differentially weight models $j$ over the decision space of $\mathbf{x},$ incorporating any prior information relevant to model weighting based on past predictive model fit and decision outcomes, and now explicitly allowing for adjustments based on a currently considered decision $\mathbf{x}$. Dependence of $\pi_{j}(\mathbf{x})$ on $\mathbf{x}$ is simply fundamental and critical in our policy setting.
@@
The $\alpha_{j}(\mathbf{y}|\mathbf{x})$ are calibration functions that define outcome-dependence of model weights over the outcome space of $\mathbf{y}$ for any chosen $\mathbf{x}$. This defines the opportunity to increase or decrease the model weights differentially over the outcome $\mathbf{y}$ space to address model-specific biases and preferences and address questions of model-specific calibration more generally. The BPDS mixture of eqn. (1) has the equivalent form
@@
Note that the choice of relevant calibration functions $\alpha_{j}(\mathbf{y}|\mathbf{x})$ will, in any given application, be partly dependent on characteristics of the model pdfs $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$. In particular, the expectation of each $\alpha_{j}(\mathbf{y}|\mathbf{x})$ under $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ must be finite in order that eqn. (1) defines a valid BPDS density $f(\mathbf{y}|\mathbf{x}).$ Unbounded score functions may sometimes apply, but this point supports the use of bounded scores in general.
@@
The model index $j=0$ explicitly allows for a baseline model component ${\mathcal{M}}_{0}$ in the mixture pdf $f(\cdot|\cdot)$ that can, among other things, address the ever-present issue of “model set incompleteness” (Tallman and West, 2023, section 2.2.3). ${\mathcal{M}}_{0}$ can be chosen to produce a pdf $f_{0}(\cdot|\cdot)$ that is over-dispersed relative to the mixture of the initial $J$ models, so supporting outcomes $\mathbf{y}$ that are unusual under the $J$ models; the baseline is then a suitable “fall back” model for times when the other models are forecasting poorly.
@@
The special case with each $\alpha_{j}(\mathbf{y}|\mathbf{x})=1$ defines the initial mixture with no BPDS calibration. We use $p(\mathbf{y}|\mathbf{x})$ in notion, i.e., $p(\mathbf{y}|\mathbf{x})=\sum_{j=1{:}J}\pi_{j}(\mathbf{x})p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$
Special cases fix ideas. First, if $\pi_{j}(\mathbf{x})=\pi_{j}$ with $\pi_{0}=0$ are model probabilities based on historical BMA analysis, and with $\alpha_{j}(\mathbf{y}|\mathbf{x})=1,$ then eqn. (1) specialises to BMA. Thus BMA analyses– with or without this decision dependence in model-specific forecasts– are very special cases of BPDS. Second, again with $\pi_{j}(\mathbf{x})=\pi_{j}$, $\pi_{0}=0$ and $\alpha_{j}(\mathbf{y}|\mathbf{x})=1,$ the decision maker has the freedom to specify the initial mixture probabilities $\pi_{j}$ in other ways than with BMA. This includes using historical performance defined by scoring of past forecast outcomes, justifying various approaches to goal-focused model weighting (e.g. Lavine et al., 2021; Loaiza-Maya et al., 2021, and references therein) as special cases of BPDS. Third, mixture BPS (McAlinn and West, 2019; Johnson and West, 2022) is a special case in which models are combined with outcome-dependent weights. In these settings, $\pi_{j}(\mathbf{x})=\pi_{j}$ depends on past predictive performance, $p_{j}(\mathbf{y}|\mathbf{x})=p_{j}(\mathbf{y})$ and $\alpha_{j}(\mathbf{y}|\mathbf{x})=\alpha_{j}(\mathbf{y})$ define outcome-dependent modifications of model probabilities, but there is no decision context so no $\mathbf{x}-$dependence. BPDS critically recognizes that the foundational BPS theory allows explicit incorporation of decision goals– admitting the conditioning on $\mathbf{x}$ throughout all components of eqn. (1)– to extend the foregoing analyses.
With predictions of $(\mathbf{y}|\mathbf{x})$ based on eqn. (2), the Bayesian decision maker acts to identify the optimal decision $\mathbf{x}$ based on a chosen utility function $U(\mathbf{y},\mathbf{x}).$ This involves numerical optimization to maximize the implied expected utility $\bar{U}(\mathbf{x})=E_{f}[U(\mathbf{y},\mathbf{x})|\mathbf{x}]$ over the decision space of $\mathbf{x}.$ The notation $E_{f}[\cdot|\cdot]$ here explicitly represents expectation with respect to the BPDS distribution, and we use $E_{p}[\cdot|\cdot]$ to denote expectation under the initial mixture.
@@
The key step in integrating decision outcomes into relative model weightings is to address the question of how each ${\mathcal{M}}_{j}$ would inform decisions if used alone. Given the predictive pdf $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ and a chosen, potentially model-specific utility function $u_{j}(\mathbf{y},\mathbf{x}),$ acting based only on ${\mathcal{M}}_{j}$ leads to the optimal decision $\mathbf{x}_{j}$ that maximizes $E_{p_{j}}[u_{j}(\mathbf{y},\mathbf{x})|\mathbf{x}]$ over $\mathbf{x}.$ The decision maker has access to this set of model recommendations and is interested in model combination to preferentially weight “good decision models” as well as models that generate good predictions. BPDS formalizes this with specified score functions $\mathbf{s}_{j}(\mathbf{y},\mathbf{x}_{j}),$ each being a $k-$vector of utilities that can be chosen to reflect both predictive and decision goals. The use of multi-dimensional scores addresses multiple goals simultaneously.
@@
BPDS takes the view that the initial mixture is based on past performance, and additional small changes in models and the way they are weighted based on their expected performance may lead to better future decisions. It asks the question as to whether there are perturbations of the mixture based on the initial model probabilities that can lead to improved scores. Consider a stylized example with $J=2$ models which in the past have forecast equally well. Traditional model averaging methods focused only on past forecasting experience– and BMA in particular– would confer equal weights in the combination. If, however, the models have different expected scores $\mathbf{m}_{jp}(\mathbf{x})$, conferring slightly more weight on the model expected to lead to a higher score makes sense. The entropic (or exponential) tilting theory is general. It is, of course, possible to tilt the initial joint distribution to most targets, so long as they are technically achievable under the initial distribution; but an overly ambitious target score will result in a tilted joint distribution that is empirically unreasonable. Hence we emphasize the importance of selecting $\mathbf{m}_{f}(\mathbf{x})$ that represents a “small ” improvement over the initial benchmark score $\mathbf{m}_{p}(\mathbf{x}).$ This is bolstered by the assumption that the initial model probabilities reflect the empirical plausibility of models as well as any available information about historical predictive and decision performance. Further, as we exemplify in the case study later, aspects of the computational methodology for model fitting in the sequential time series setting naturally inform on, and allow monitoring of, relevant choices of target expected scores.
@@
This section has outlined the main ideas underlying BPDS and the key ingredients of the theory and resulting technical machinery. Specifications of score and utility functions, initial model probabilities, and target scores are all required for implementation and are, of course, application specific. The following section develops full details in the context of the macroeconomic decision-making application. In terms of computation, BPDS requires the use of posterior simulation methods (i.e. draws from conditional predictive densities from each model are required) as well as numerical optimization methods (i.e. to find $\mathbf{x}_{j}$ or the overall optimal decision $\mathbf{x}$ under the final BPDS analysis).
@@
The choice of data and models is inspired by Furlanetto et al. (2019). We use quarterly macroeconomic and financial variables from 1973:Q1 to 2022:Q2 from the FRED-QD database maintained by the Federal Reserve Bank of St. Louis. The data set includes GDP (the log of real GDP), prices (the log of the GDP deflator), the interest rate (the shadow rate [^1] This is the Federal Funds rate when the latter is positive but can go negative when it is at the zero lower bound, taking into account unconventional monetary policy; see Wu and Xia (2016). which we treat as the policy rate), investment (the ratio of real gross private domestic investment to GDP), stock prices (log of the S&P500) and the spread (the spread between BAA bonds and the Fed funds rate). Models are run over multiple years, and at the end of each quarter produce forecasts– full predictive distributions in terms of Monte Carlo samples– of outcomes of interest over the following $k=8$ quarters, conditional on candidate settings of the decision vector which is taken as the trajectory of interest (shadow) rates over those quarters. Within-model decision analysis then delivers model-specific optimal decisions about these rates.
Models, Forecasts and Model-Specific Decisions
@@
We consider $J=2$ models: ${\mathcal{M}}_{1}$ is a three-variable monetary policy VAR involving GDP, prices, and the interest rate; ${\mathcal{M}}_{2}$ is the model of Furlanetto et al. (2019), a VAR with the same variables as ${\mathcal{M}}_{1}$ plus investment, stock prices, and the spread. Following the latter, we include 5 lags in the VARs. The two structural VARs are identified using the sign restrictions from Table 1 of Furlanetto et al. (2019). In ${\mathcal{M}}_{1}$ these restrictions define supply, demand and monetary policy shocks. In ${\mathcal{M}}_{2}$ investment and financial shocks are additionally identified. We condition on a given value of the policy rate and set the monetary policy to be the driving shock. We do this by imposing restrictions on the set of structural shocks underlying the conditional forecasts. Structural shocks other than the monetary policy shock have zero means. We use the asymmetric conjugate prior of Chan (2022), with the advantage that the marginal likelihoods for each can be easily calculated; prior hyperparameter choices are made to maximize the marginal likelihood as in this referenced paper. At each quarter, multi-step ahead predictions are based on simulations using the precision-based sampler of Chan et al. (2023). Details on the conditional forecast computations are summarized in Appendix B. In short, this generates $p_{j}(\mathbf{y},|\mathbf{x})$ with zero-mean constraints on all shocks apart from the monetary policy variable. We do not restrict the variance (i.e., ‘soft’ restrictions) such that we also have uncertainty around the path of $\mathbf{x}$. This can be thought of as conditional commitment– we allow the possibility of $\mathbf{x}$ deviating from the proposed policy path with deviations informed by historical uncertainty around forecasts of the interest rate outcomes.

### v2

@@ Abstract
The macroeconomy is a sophisticated dynamic system involving significant uncertainties that complicate modelling. In response, decision-makers consider multiple models that provide different predictions and policy recommendations which are then synthesized into a policy decision. In this setting, we develop Bayesian predictive decision synthesis (BPDS) to formalize monetary policy decision processes. BPDS draws on recent developments in model combination and statistical decision theory that yield new opportunities in combining multiple models, emphasizing the integration of decision goals, expectations and outcomes into the model synthesis process. Our case study concerns central bank policy decisions about target interest rates with a focus on implications for multi-step macroeconomic forecasting. This application also motivates new methodological developments in conditional forecasting and BPDS, presented and developed here.
@@ Introduction
Monetary policymakers are tasked with simple, but hard to achieve, objectives. A key example is the “dual mandate” under which policymakers target future inflation rates and real activity using interest rates as the policy instrument. Decisions are made based on uncertain information from many sources. Here, the sources are models that generate predictive distributions for macroeconomic outcomes and policy instruments over multiple time periods. For a single model, an optimal policy path is evaluated via conditional forecasting and decision analysis. Applying standard methods, such as Bayesian model averaging (BMA), is one way to address model uncertainty– routine decision analysis can then be applied to the weighted average of models. This traditional view, however, ignores the reality that models may each individually recommend very different optimal policy decisions. The question then arises of how to synthesize this information and, potentially, exploit it in the overall final decision process. This paper addresses this question.
@@
The extensive Bayesian econometrics literature on model combination rarely highlights the fact that models are typically built for specific prediction and decision goals. Traditional BMA analysis weights models according to purely statistical model fit and only scores one-step ahead forecast outcomes. Extensions and alternatives have arisen to define model weightings based on aspects of past forecast performance with respect to specific forecast goals. Martin et al. (2023) survey Bayesian forecasting in economics and finance and review various forecast combination approaches, including some that are more explicitly concerned with goal-focused prediction (e.g., Mitchell and Hall, 2005; Geweke and Amisano, 2011; Conflitti et al., 2015; Kapetanios et al., 2015; Loaiza-Maya et al., 2021; Chernis and Webley, 2022; Aastveit et al., 2023; Bernaciak and Griffin, 2024). Lavine et al. (2021) place many of the earlier approaches in a foundational Bayesian context. They justify model weights based on utilities in forecasting using historical model-specific “scoring” of past forecast outcomes. The underlying theoretical justifications come from Bayesian predictive synthesis (BPS) and the specific class of “mixture BPS” models (McAlinn and West, 2019, section 2.2; Johnson and West, 2025). However, while ultimate decision goals may be implicit in specific applications of model combination, they are rarely, if ever, taken into account in the analysis and resulting decision-making. This raises concerns. A model that has fit or forecast specific outcomes well in the past may be a good bet for use in resulting decision analysis, such as defining optimal decisions about values of policy instruments; however, there is no guarantee that this will be so.
@@
Our view is that models that have recommended policy decisions that turned out to be “good” should be more heavily weighted in looking ahead, just as past statistical predictive performance is generally positively weighted. The challenge is to operationalize the concept of “good decision” performance. For example, a vector autoregression (VAR) model can be evaluated on forecast performance using a pseudo real-time forecasting exercise, but it is not clear how to evaluate such a model when it is used to advise policy decisions. We can, however, explore how the model would have advised on decisions in the past. Evaluations can then compare such analyses to decisions actually made by policymakers in the past (while recognizing that policymakers’ past decisions were not necessarily “good” and just outcomes of the rather amorphous reality of monetary policymaking).
@@
Bayesian predictive decision synthesis (BPDS– Tallman and West, 2023) has the potential to address these questions. An outgrowth of the theoretical BPS framework, BPDS explicitly addresses model scoring based on decision outcomes as well as predictive accuracy. In addition to reflecting historical outcomes of predictions and decisions, BPDS allows for differential model weighting based on expected decision outcomes. This is a decision parallel to the proven use of BPS models that incorporate outcome-dependent weights that modify BMA-like mixtures to differentially favour models in different parts of the future outcome space for pure forecasting. The latter concept was introduced by Kapetanios et al. (2015) whose empirically inspired developments recognized, for example, that one model may be better at predicting inflation when inflation is high and rising, while another model may be better when inflation is low and stable. BPS defines a theoretical Bayesian basis and broader methodological framework for this (Johnson and West, 2025). BPDS goes further by integrating both historical and expected decision outcomes; here we develop, extend, and exemplify BPDS in the central macroeconomic policy context.
@@
BPDS applies the Bayesian mixture model approach of BPS using defined utility– or “score”– functions that relate to explicit decision goals. This allows for multiple objectives (i.e., multi-attribute decision analysis). For example, a purely predictive vector score function can allow for multiple forecast horizons (e.g., to produce inflation near a target for each of the next eight quarters) and/or multiple outcome criteria (e.g., to separately reflect inflation targeting, interest rate smoothing, and stable growth patterns over coming quarters), among others. For policymakers juggling multiple objectives, this key feature is rather distinct from conventional approaches that adopt single, scalar criteria for model weighting. For example, a forecast combination approach might choose model weights based on the $h-$step ahead predictive likelihood for a single choice of $h$, with BMA simply focused on $h=1$. In contrast, BPDS can address multi-steps ahead in parallel, along with scoring of realized decisions that simultaneously target several macroeconomic outcomes.
The opportunities for exploring such practically relevant questions are highlighted in empirical studies. Our case study uses US macroeconomic data with multi-objective score functions to define BPDS model weights, demonstrating BPDS in macroeconomic forecasting and advisory decision-making. This is enabled by new methodological advances motivated by this applied context. First, in developing the scenario-based conditional forecasting analyses, we explicitly recognize the dual roles of policy decision variables (such as interest rates) as both outcomes and putative controls. This leads to an important theoretical extension of traditional conditional forecasting in which candidate decision scenarios are weighted by their predicted plausibility prior to assessing their likely implications in forecasting outcomes of other variables. [^1] This is keeping in the spirit of Leeper and Zha (2003) who argue that the “modest interventions”commonly used in reduced-form scenario analysis are unlikely to be subject to the Lucas Critique. We explicitly give higher weight to small interventions and down-weight large interventions. This is important quite generally in conditional forecasting, as well as here in the resulting BPDS analysis. Second, we advance the BPDS methodology based on a new and practically critical focus on the relationships between multiple utilities defining the scoring of decision outcomes. Third, we discuss and promote the use of bounded utilities in model scoring, and customized forms of utility functions to guide model-specific and final BPDS-based optimal decisions.
@@
At a given time point, let $\mathbf{y}$ denote the $q-$dimensional outcome variable of interest (e.g., inflation in each of the next $q$ quarters) and $\mathbf{x}$ the vector of control/decision variables (e.g., a target profile of central bank interest/base rates over the next $q$ quarters). Each of a set of $J$ models, ${\mathcal{M}}_{j}$, $j=1{:}J,$ predicts the outcome $\mathbf{y}$ via a predictive density $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ conditional on any considered decision $\mathbf{x}.$ The policymaker responsible for ultimate decisions adopts a general BPDS approach with the overall conditional (on $\mathbf{x}$) predictive p.d.f.
@@
The decision-dependent model probabilities $\pi_{j}(\mathbf{x})$ can differentially weight models $j$ over the decision space $\mathbf{x}$. This incorporates any prior information relevant to model weighting based on past predictive model fit and decision outcomes, and now explicitly allows for adjustments based on a currently considered decision $\mathbf{x}$. Dependence of $\pi_{j}(\mathbf{x})$ on $\mathbf{x}$ is simply fundamental and critical in our policy setting.
@@
The $\alpha_{j}(\mathbf{y}|\mathbf{x})$ are calibration functions that define outcome dependence of model weights over the outcome space of $\mathbf{y}$ for any chosen $\mathbf{x}$. This defines the opportunity to increase or decrease the model weights differentially over the outcome $\mathbf{y}$ space to address model-specific biases and preferences and address questions of model-specific calibration more generally. The BPDS mixture of eqn. (1) has the equivalent form
@@
Note that the choice of relevant calibration functions $\alpha_{j}(\mathbf{y}|\mathbf{x})$ will, in any given application, be partly dependent on characteristics of the model pdfs $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$. In particular, the expectation of each $\alpha_{j}(\mathbf{y}|\mathbf{x})$ under $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ must be finite in order that eqn. (1) defines a valid BPDS density $f(\mathbf{y}|\mathbf{x}).$ This supports the use of bounded scores, in general.
@@
The model index $j=0$ explicitly allows for a baseline model component ${\mathcal{M}}_{0}$ in the mixture p.d.f. $f(\cdot|\cdot)$ that can, among other things, address the ever-present issue of “model set incompleteness” (Tallman and West, 2023, section 2.2.3). ${\mathcal{M}}_{0}$ can be chosen to produce a pdf $f_{0}(\cdot|\cdot)$ that is over-dispersed relative to the mixture of the initial $J$ models, so supporting outcomes $\mathbf{y}$ that are unusual under the $J$ models. The baseline is then a suitable “fall back” model for times when the other models are forecasting poorly.
@@
The special case with each $\alpha_{j}(\mathbf{y}|\mathbf{x})=1$ defines the initial mixture with no BPDS calibration. We use $p(\mathbf{y}|\mathbf{x})$ in notion; that is, $p(\mathbf{y}|\mathbf{x})=\sum_{j=1{:}J}\pi_{j}(\mathbf{x})p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$.
Special cases fix ideas. First, if $\pi_{j}(\mathbf{x})=\pi_{j}$ with $\pi_{0}=0$ are model probabilities based on historical BMA analysis, and with $\alpha_{j}(\mathbf{y}|\mathbf{x})=1,$ then eqn. (1) specializes to BMA. Thus, BMA analyses– with or without this decision dependence in model-specific forecasts– are very special cases of BPDS. Second, again with $\pi_{j}(\mathbf{x})=\pi_{j}$, $\pi_{0}=0$ and $\alpha_{j}(\mathbf{y}|\mathbf{x})=1,$ the decision-maker has the freedom to specify the initial mixture probabilities $\pi_{j}$ in other ways than with BMA. This includes using historical performance defined by scoring of past forecast outcomes, justifying various approaches to goal-focused model weighting (e.g., Lavine et al., 2021; Loaiza-Maya et al., 2021, and references therein) as special cases of BPDS. Third, mixture BPS (McAlinn and West, 2019; Johnson and West, 2025) is a special case in which models are combined with outcome-dependent weights. In these settings, $\pi_{j}(\mathbf{x})=\pi_{j}$ depends on past predictive performance, $p_{j}(\mathbf{y}|\mathbf{x})=p_{j}(\mathbf{y})$ and $\alpha_{j}(\mathbf{y}|\mathbf{x})=\alpha_{j}(\mathbf{y})$ define outcome-dependent modifications of model probabilities, but there is no decision context so no $\mathbf{x}-$dependence. BPDS critically recognizes that the foundational BPS theory allows explicit incorporation of decision goals– admitting the conditioning on $\mathbf{x}$ throughout all components of eqn. (1)– to extend the foregoing analyses.
With predictions of $(\mathbf{y}|\mathbf{x})$ based on eqn. (2), the Bayesian decision-maker acts to identify the optimal decision $\mathbf{x}$ based on a chosen utility function $U(\mathbf{y},\mathbf{x}).$ This involves numerical optimization to maximize the implied expected utility $\bar{U}(\mathbf{x})=E_{f}[U(\mathbf{y},\mathbf{x})|\mathbf{x}]$ over the decision space of $\mathbf{x}.$ The notation $E_{f}[\cdot|\cdot]$ here explicitly represents expectation with respect to the BPDS distribution, and we use $E_{p}[\cdot|\cdot]$ to denote expectation under the initial mixture.
@@
The key step in integrating decision outcomes into relative model weights addresses the question of how each ${\mathcal{M}}_{j}$ would inform decisions if used alone. Given the predictive p.d.f. $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ and a chosen, potentially model-specific utility function $u_{j}(\mathbf{y},\mathbf{x}),$ acting based only on ${\mathcal{M}}_{j}$ leads to the optimal decision $\mathbf{x}_{j}$ that maximizes $E_{p_{j}}[u_{j}(\mathbf{y},\mathbf{x})|\mathbf{x}]$ over $\mathbf{x}.$ The decision-maker has access to this set of model recommendations and is interested in model combination to preferentially weight “good decision models” as well as models that generate good predictions. BPDS formalizes this with specified score functions $\mathbf{s}_{j}(\mathbf{y},\mathbf{x}_{j}),$ each being a $k-$vector of utilities that can be chosen to reflect both predictive and decision goals. The use of multi-dimensional scores addresses multiple goals simultaneously.
@@
BPDS uses the initial mixture based on past performance, but then admits that small perturbations of model probabilities based on their expected performance may lead to improved scores. A stylized example has $J=2$ models that have forecast equally well to date. Traditional model averaging methods including BMA confer equal weights in the combination. If, however, the models have different expected scores $\mathbf{m}_{jp}(\mathbf{x})$, conferring slightly more weight on the model expected to lead to a higher score makes sense.
The entropic (or exponential) tilting theory is general. It is, of course, possible to tilt the initial joint distribution to most targets, so long as they are technically achievable under the initial distribution. However, an overly ambitious target score may result in empirically unreasonable results. Hence, we emphasize the importance of selecting $\mathbf{m}_{f}(\mathbf{x})$ that represents a “small” improvement over the initial benchmark score $\mathbf{m}_{p}(\mathbf{x}).$ This is bolstered by the assumption that the initial model probabilities reflect the empirical plausibility of models, as well as any available information about historical predictive and decision performance. Further, as we demonstrate later in the case study, aspects of the computational methodology for model fitting in the sequential time series setting naturally inform on, and allow monitoring of, relevant choices of target expected scores.
@@
This section has outlined the main ideas underlying BPDS and the key ingredients of the theory and resulting technical machinery. Specifications of score and utility functions, initial model probabilities, and target scores are all required for implementation and are, of course, application specific. The following section develops full details in the context of the macroeconomic decision-making application. In terms of computation, BPDS requires the use of posterior simulation methods (i.e., draws from conditional predictive densities from each model), as well as numerical optimization methods (i.e., to find the $\mathbf{x}_{j}$ and the overall optimal decision $\mathbf{x}$ from BPDS analysis).
@@
Following Furlanetto et al. (2019), we use quarterly macroeconomic and financial data from 1973:Q1 to 2022:Q2 from the FRED-QD database (Federal Reserve Bank of St. Louis). This includes GDP (log of real GDP), prices (log of GDP deflator), interest rate (the shadow rate, [^2] This is the Federal Funds rate when the latter is positive, but can go negative when it is at the zero lower bound, taking into account unconventional monetary policy; see Wu and Xia (2016). which we treat as the policy rate), investment (ratio of real gross private domestic investment to GDP), real stock prices (log of the S&P500 deflated by GNP/GDP price index), and spread (between BAA bonds and the Fed funds rate). Models are run over multiple years. Each quarter they produce forecasts– full predictive distributions in terms of Monte Carlo samples– of outcomes of interest over the following $k=8$ quarters. This is conditional on candidate settings of the decision vector, which is taken as the trajectory of interest (shadow) rates over those quarters. Within-model decision analysis then delivers model-specific optimal decisions about these rates.
Models, Forecasts, and Model-specific Decisions
@@
We consider $J=2$ models: ${\mathcal{M}}_{1}$ is a three-variable monetary policy VAR involving GDP, prices, and the interest rate; ${\mathcal{M}}_{2}$ is similar to the model of Furlanetto et al. (2019), a VAR with the same variables as ${\mathcal{M}}_{1}$ plus investment/GDP ratio, stock prices, and GZ credit spread. Following Furlanetto et al. (2019) we include five lags in the VARs, and each model is identified using the sign restrictions from Table 1 of that reference. In ${\mathcal{M}}_{1}$, these restrictions define supply, demand, and monetary policy shocks. In ${\mathcal{M}}_{2}$, investment and financial shocks are additionally identified. We condition on a given value of the policy rate and set monetary policy to be the driving shock. We do this by imposing restrictions on the set of structural shocks underlying the conditional forecasts. Structural shocks other than the monetary policy shock have zero means. We use the asymmetric conjugate prior of Chan (2022), with the advantage that the marginal likelihoods for each can be easily calculated; prior hyperparameter choices are made to maximize the marginal likelihood as in this referenced paper. At each quarter, multi-step ahead predictions are based on simulations using the precision-based sampler of Chan et al. (2025). Details on the conditional forecast computations are summarized in Appendix B. In short, this generates $p_{j}(\mathbf{y}|\mathbf{x})$ with zero-mean constraints on all shocks apart from the monetary policy variable. We do not restrict the variance (i.e., “soft” restrictions) such that we also have uncertainty around the path of $\mathbf{x}$. This can be thought of as conditional commitment– we allow the possibility of $\mathbf{x}$ deviating from the proposed policy path with deviations informed by historical uncertainty around forecasts of the interest rate outcomes.
It is worth noting that using models such as VARs for policymaking is sometimes questioned due to the Lucas critique. However, papers such as Leeper and Zha (2003) argue in favor of VARs as long as the policy interventions being studied are modest. We ensure BPDS will favor such modest interventions through two mechanisms. First, as noted in Section 2.2, we choose only modest improvements in the target score. Second, large policy interventions will be down-weighted through the initial probabilities reflecting the plausibility of an intervention; see details in Section 3.2. Thus BDPS allows for policy analysis using VARs by emphasizing decision scenarios which represent “modest interventions”, circumventing the Lucas Critique. This is an important practical contribution to the VAR literature on policy and scenario analysis.